# Nonlinearity-induced Thouless pumping of solitons

### v1

@@ Abstract
To further characterize the topological property of anomalous nonlinear Thouless pumping, we calculate the Chern number of a modulated linear Hamiltonian in a supercell consisting of $L$ unit cells incorporating the effects of a soliton [34] (see also Supplementary Information Section S-3). We choose an $L$ (here we take $L=40$) such that the soliton is well localized within the supercell. As we vary $\theta$ from 00 to $2L\pi$, the soliton is pumped across $Ld_{p}$ unit cells where $d_{p}$ is the displacement over each cycle of $\theta$ from 00 to $2\pi$. When the soliton travels outside the supercell, we shift it by $L$ back into the supercell. By involving the effects of the soliton, we define a new linear Hamiltonian $H^{\textrm{sc}}(\theta)$ as
$H^{\textrm{sc}}_{\sigma x,\sigma^{\prime}x^{\prime}}:=H^{\textrm{lin}}_{\sigma x,\sigma^{\prime}x^{\prime}}+(g|\psi_{\sigma x}^{\textrm{s}}|^{2}+g_{12}|\psi_{\overline{\sigma}x}^{\textrm{s}}|^{2})\delta_{xx^{\prime}}\delta_{\sigma\sigma^{\prime}},$ (3)
where $\psi_{\sigma x}^{\textrm{s}}$ is the soliton solution we consider at a fixed $\theta$. Under periodic boundary conditions, $\psi_{\sigma x}^{\textrm{s}}$ is clearly an eigenstate of $H^{\textrm{sc}}$. By applying twisted boundary condition for the Hamiltonian $H^{\textrm{sc}}$ with a twisted phase $k_{\textrm{sc}}$ at each $\theta$, we calculate the Chern number $C_{\textrm{sc}}$ of the correponding band of $H^{\textrm{sc}}$ with respect to $\theta^{\prime}=\theta/L$ and $k_{\textrm{sc}}$. We find that $C_{\textrm{sc}}=d_{p}$ when $d_{p}=0,1,2$ in our cases. The results are self-consistent as the soliton in a supercell plays the role of the Wannier function of $H^{\textrm{sc}}$ consisting of many supercells (our numerical results also support this viewpoint). As a result, the soliton will travel across $C_{\textrm{sc}}$ supercells after a period of $\theta^{\prime}$ so that the displacement over one cycle of $\theta$ is $C_{\textrm{sc}}$. The anomalous nonlinear Thouless pumping is thus incorporated into the framework of the supercell theory.
@@
Continuous model.— To implement the tight-binding model in ultracold atomic gases, we consider the following continuous model:
$H_{\textrm{c}}^{\textrm{lin}}=H_{0}+h_{z}\sigma_{z}+V_{\textrm{so}}\sigma_{x},$ (4)
@@
where the spin is encoded in two hyperfine states of an atom. Note that similar models have been proposed for realizing linear Thouless pumping in both two-dimensional and three-dimensional cold atom systems [42]. In this Hamiltonian, $H_{0}=\frac{p_{x}^{2}}{2m}-V_{x}\cos^{2}(k_{R}x)$, where $p_{x}=-i\hbar\partial_{x}$ is the momentum operator, $m$ is the mass of atoms, and $V_{x}$ is the strength of optical lattices generated by lasers with wavevector $k_{R}$. In addition, $h_{z}$ denotes the strength of the Zeeman field correponding to the detuning of the Raman lasers, and $V_{\textrm{so}}=V_{s}\sin(k_{R}x)+V_{c}\cos(k_{R}x)\sin(\theta)$. This potential can be implemented using two pairs of Raman lasers, as described in Ref. [42, 48, 49, 50]. One pair (the other) contains a laser beam with the Rabi frequency proportional to $\sin(k_{R}x)$ [$\cos(k_{R}x)$] and another beam proportional to $e^{ik_{R}y}$. Given that our system is effectively one-dimensional due to a strong confinement along $y$ and $z$, the effects of the phase $e^{ik_{R}y}$ can be negligible. We can map the continuous model to a tight-binding model based on the Wannier functions $\{W_{j}\}$ of the lowest Bloch band of the Hamiltonian $H_{0}$, yielding precisely the Hamiltonian in Eq. (2) with $J_{1}=2t_{1}$, $J_{1}^{\prime}=-2t_{s}$, $J_{2}=t_{c}\sin\theta$ and $m_{z}=h_{z}$ (see Supplementary Information Section S-4 for their definition). Thus, we expect the emergence of anomalous nonlinear Thouless pumping in the continuous model in the presence of interactions described by nonlinear terms.
@@
For the linear Thouless pumping of this model, we prove that the average displacement per particle at time $t$ is $x_{c}=a[\gamma(\theta(t))-\gamma(\theta(t=0))]/(2\pi)$ where $\gamma(\theta(t))$ is the Berry phase following the path along $k$ at a fixed $\theta$ and $a=\pi/k_{R}$, although the length of a unit cell of $H_{c}^{\textrm{lin}}$ is $2a$ due to the presence of $V_{\textrm{so}}$ (see Supplementary Information Section S-5 for proof). Here we require that $\gamma(\theta(t))$ changes continuously from $\theta(0)$ to $\theta(t)$. In fact, we find that the linear eigenstates can be characterized by the momentum $k\in[0,2\pi/a]$. In our study, we set $V_{x}=4E_{R}$, $V_{s}=1.04E_{R}$ and $V_{c}=0.21E_{R}$, where $E_{R}=\hbar^{2}k_{R}^{2}/(2m)$ is the recoil energy. The parameter $h_{z}$ is varied as $h_{z}=2t_{x}(1+\cos\theta)$ with $t_{x}=0.0855E_{R}$. Under these conditions, we find that the Chern number of the first and second bands of the continuous model are $C=1$ and $C=-1$, respectively. As a result, the average displacement for the linear Thouless pumping over one cycle is $a$ and $-a$, respectively (see Fig. 4b).
Figure 4: Anomalous nonlinear pumping in the continuous model. a, One-cycle evolution of the density distribution $|\psi(x)|^{2}=\sum_{\sigma}|\psi_{\sigma}(x)|^{2}$ of the instantaneous solitons bifurcating from the second band when $g_{12}=0$. b, The evolution of center-of-mass positions $x_{c}$ of the instantaneous solitons (red line) and the corresponding Wannier functions (black line) with respect to $\theta$ over one period. Here, ${N}=0.2$ and $g=-1$.
@@
We now demonstrate the emergence of anomalous nonlinear pumping in the continuous model in the presence of nonlinearity. The dynamics is governed by the following dimensionless Gross-Pitaevskii equation:
@@
[…]
In addition, we show that for negative $g$, a soliton bifurcating from the first band does not exhibit any pumping over one cycle. Although the phenomenon is analogous to a trapped soliton [34], the underlying mechanism differs markedly. The trapped phenomenon arises from multiple band contributions due to strong nonlinearity [34, 36, 40]. In contrast, in our case, strong nonlinearity is not a prerequisite, and the projections of the soliton indicate primary occupation of a single band. This behavior in fact stems from the emergence of a new branch of soliton solutions, which is primarily composed of two adjacent instantaneous Wannier functions (see Supplementary Information Section S-6).
@@
Discussion.— We have demonstrated the emergence of anomalous nonlinear Thouless pumping, wherein the correspondence between the displacement of a soliton and the Chern number of the underlying linear Bloch band breaks down. This anomalous behavior surpasses the understanding of nonlinear Thouless pumping based on the center-of-mass flow of the instantaneous Wannier functions. We show that the breakdown arises due to the emergence of a new branch of soliton solutions primarily composed of two neighboring instantaneous Wannier functions. Furthermore, we illustrate that nonlinearity can induce the Thouless pumping of a soliton, even when the corresponding linear Bloch band is topologically trivial. We note that our results are completely different from the phenomenon in a Bose-Bose atomic mixture in Ref. [37] where a soliton of impurity atoms can undergo a nonzero quantized displacement during a pump cycle, as a result of the dragging from a soliton of majority atoms that bifurcates from a topologically nontrivial band. There, the pumping can also be understood as the center-of-mass flow of the Wannier functions of the corresponding Bloch band of the linear Hamiltonian felt by the majority atoms. In our case, we have only one set of band structures, and the discrepancy between the band topology and quantized motion emerges spontaneously, going beyond the Wannier function description. Given that anomalous nonlinear Thouless pumping can occur in both discrete and continuous models, it can be experimentally observed in either photonic systems or ultracold atomic gases. Additionally, our work inspires the exploration of other interesting topics, such as anomalous fractional Thouless pumping of solitons.

### v2

@@ Abstract
To characterize the topological property of anomalous nonlinear Thouless pumping, we calculate the Chern number of a modulated linear Hamiltonian in a supercell consisting of $L$ unit cells incorporating the effects of solitons [35]. To establish the validity of this method, consider a soliton solution $\psi_{\sigma j}^{\textrm{s}}(\theta)$ that satisfies Eq. (1). This solution is used to define a new linear Hamiltonian at each $\theta$ as
$H^{\prime}_{\sigma j,\sigma^{\prime}j^{\prime}}(\theta):=H^{\textrm{lin}}_{\sigma j,\sigma^{\prime}j^{\prime}}(\theta)+V_{\sigma j}(\psi_{\sigma j}^{\textrm{s}}(\theta),\psi_{\overline{\sigma}j}^{\textrm{s}}(\theta))\delta_{jj^{\prime}}\delta_{\sigma\sigma^{\prime}},$ (3)
where $V_{\sigma j}(\psi_{\sigma j}^{\textrm{s}}(\theta),\psi_{\overline{\sigma}j}^{\textrm{s}}(\theta))$ is the effective potential generated by the soliton. We then study the time-dependent linear equation, $i\partial_{t}\Psi_{\sigma j}=\sum_{\sigma^{\prime}j^{\prime}}H^{\prime}_{\sigma j,\sigma^{\prime}j^{\prime}}(\theta)\Psi_{\sigma^{\prime}j^{\prime}}$, where $H^{\prime}$ is independent of $\Psi$. Clearly, $\psi_{\sigma j}^{\textrm{s}}(\theta)$ is a solution to this linear equation, thus converting a nonlinear problem to a linear one. However, $H^{\prime}$ does not preserve translational symmetry due to the soliton-induced defect potential. To restore the symmetry, we partition the lattice into sections $\dots,[1,L],[L,2L],\dots$, each called a supercell, composed of $L$ unit cells. $L$ is sufficiently large so that a soliton is well localized within a supercell at $\theta=0$. At this $\theta$, a soliton exists within each supercell, as illustrated in Fig. 5a, making the effective potential caused by the solitons periodic with a period of $L$. Since the solitons are far away from each other, the soliton array should also be the solution to Eq. (1). Similar to $H^{\prime}(\theta)$, we define a linear Hamiltonian $H^{\textrm{sc}}(\theta)$ using the effective potential of the soliton array at each $\theta$, which respects the translational symmetry, i.e., $H^{\textrm{sc}}_{\sigma j+L,\sigma^{\prime}j^{\prime}+L}(\theta)=H^{\textrm{sc}}_{\sigma j,\sigma^{\prime}j^{\prime}}(\theta)$.
Our numerical calculations indicate that the soliton in each supercell closely resembles a Wannier function of the Hamiltonian $H^{\textrm{sc}}(\theta)$. This is evidenced by the uniform projections of the soliton wavefunction onto the second band in Fig. 5b, as shown in Fig. 5c, indicating that the soliton wavefunctions are equal superpositions of all the Bloch states in this band, and are thus Wannier functions. Since the displacement of a Wannier function over a pump period is determined by the Chern number $C_{\textrm{sc}}$ of this band of $H^{\textrm{sc}}$, the soliton’s displacement is thus dictated by this Chern number. To see this, we introduce a new parameter $\theta^{\prime}=\theta/L$ such that $H^{\textrm{sc}}(\theta^{\prime}+2\pi)=H^{\textrm{sc}}(\theta^{\prime})$. As $\theta^{\prime}$ varies from 0 to $2\pi$, the Wannier function travels by $C_{\textrm{sc}}$ supercells, equivalent to $C_{\textrm{sc}}L$ original unit cells, indicating that the average displacement over a cycle of $\theta$ is $C_{\textrm{sc}}$. Our calculations reveal that for normal nonlinear pumping, $C_{\textrm{sc}}$ is equal to the Chern number $C$ of the corresponding Bloch band of $H^{\textrm{lin}}$; however, they differ in the anomalous one. For example, in the anomalous cases shown in Fig. 2, $C_{\textrm{sc}}$ can be 0, 2-2, or 3-3 , while C=1C=-1. In Fig. 4, $C_{\textrm{sc}}=-1$ while $C=0$. In addition, we find that $C_{\textrm{sc}}$ changes from 1-1 to 2-2 as $g_{12}$ decreases across $0.2$ in Fig. 3.
Figure 5: Supercell method. a, The effective potential generated by a soliton array with one soliton per supercell (three supercells are explicitly shown, each containing $L=40$ sites) as $\theta^{\prime}=\theta/L$ varies from 0 to $2\pi$. The potential is periodic with period $L$. Only the potential felt by the first component is plotted, and the potential for the other component is similar. b, Energy spectra of the modulated linear Hamiltonian $H^{\textrm{sc}}$ containing the linear potentials in a with respect to momentum $k$ and $\theta^{\prime}$. Given that the spectra are periodic functions of $\theta^{\prime}$ with a period of $2\pi/L$, we only plot part of the spectrum with $\theta^{\prime}\in[0,2\pi/L]$. Here, we show three Bloch bands around $\mu(\theta)$ and the second band corresponds to the soliton wavefunction $\psi_{\sigma j}^{\textrm{s}}(\theta)$. c, The normalized projection of a soliton wavefunction $\psi_{\sigma j}^{\textrm{s}}$ on the second band in b, $\rho_{\textrm{sc}}(k,\theta^{\prime})=\frac{L_{\textrm{sc}}}{N}|\sum_{\sigma j}\varphi_{k,\sigma,j}(\theta^{\prime})^{*}\psi_{\sigma j}^{\textrm{s}}(\theta^{\prime})|^{2},$ where $L_{\textrm{sc}}$ is the number of supercells, $\varphi_{k,\sigma,j}(\theta^{\prime})$ is a Bloch state at momentum $k\in[0,2\pi/L]$ in the corresponding band of the Hamiltonian $H^{\textrm{sc}}$. Here, we consider case 2 in Fig. 2 with $N=1.45$, $m_{0}=1$, $g_{12}=0$, and $g=1$.
@@
Continuous nonlinear model.—To implement the tight-binding model in Eq. (2) in ultracold atomic gases, we consider the following continuous model:
$H_{\textrm{c}}^{\textrm{lin}}{(x,\theta)}=H_{0}{(x)}+h_{z}{(\theta)}\sigma_{z}+V_{\textrm{so}}{(x,\theta)}\sigma_{x},$ (4)
@@
where the spin is encoded in two hyperfine states of an atom. Note that similar models have been proposed for realizing linear Thouless pumping in both two-dimensional and three-dimensional cold atom systems [54]. In this Hamiltonian, $H_{0}{(x)}=\frac{p_{x}^{2}}{2m}-V_{x}\cos^{2}(k_{R}x)$, where $p_{x}=-i\hbar\partial_{x}$ is the momentum operator, $m$ is the mass of atoms, and $V_{x}$ is the strength of optical lattices generated by lasers with wavevector $k_{R}$. In addition, $h_{z}$ denotes the strength of the Zeeman field corresponding to the detuning of the Raman lasers, and $V_{\textrm{so}}{(x,\theta)}=V_{s}\sin(k_{R}x)+V_{c}\cos(k_{R}x)\sin(\theta)$. This potential can be implemented using two pairs of Raman lasers, as described in Ref. [54, 60, 61, 62]. One pair (the other) contains a laser beam with the Rabi frequency proportional to $\sin(k_{R}x)$ $\cos(k_{R}x)$] and another beam proportional to $e^{ik_{R}y}$. Given that our system is effectively 1D due to a strong confinement along $y$ and $z$, the effects of the phase $e^{ik_{R}y}$ can be negligible. The Hamiltonian $H_{\textrm{c}}^{\textrm{lin}}(x,\theta)$ is invariant under translation by $2a$ with $a=\pi/k_{R}$. Despite this, we show that its Bloch states can still be characterized by the momentum $k$ within the interval $0,2\pi/a]$, due to the symmetry represented by $T_{a}\sigma_{z}$, where $T_{a}$ is the translation operator by $a$ (see Supplementary Information Section S-6). As a result, for the linear Thouless pumping in this model, the average displacement per particle over one pump period is $x_{c}=Ca$, where $C$ is the Chern number for the corresponding Bloch band (see Supplementary Information Section S-6).
@@
The continuous model can be mapped to a tight-binding model based on the Wannier functions $\{W_{nj}(x):j\in\mathbb{Z}\}$ from the lowest Bloch band of the Hamiltonian $H_{0}(x)$ with the band index $n=1$. This yields precisely the Hamiltonian in Eq. (2) with the parameters $J_{1}=-2\int{\rm d}xW_{10}(x)H_{0}(x)W_{11}(x)$, $J_{1}^{\prime}=-2V_{s}\int{\rm d}xW_{10}(x)\sin(k_{R}x)W_{11}(x)$, $J_{2}=V_{c}\sin\theta\int{\rm d}xW_{10}(x)\cos(k_{R}x)W_{10}(x)$, and $m_{z}=h_{z}$ (see the derivation in Supplementary Information Section S-6). Thus, we expect the emergence of anomalous nonlinear Thouless pumping in the continuous model in the presence of interactions described by nonlinear terms. In our study, we set $V_{x}=4E_{R}$, $V_{s}=1.04E_{R}$ and $V_{c}=0.21E_{R}$, where $E_{R}=\hbar^{2}k_{R}^{2}/(2m)$ is the recoil energy. The parameter $h_{z}$ is varied as $h_{z}=2t_{x}(1+\cos\theta)$ with $t_{x}=0.0855E_{R}$. Under these conditions, we find that the Chern number of the first and second bands of the continuous model are $C=1$ and C=1C=-1, respectively. As a result, the average displacement for the linear Thouless pumping over one cycle is $a$ and a-a, respectively (see Fig. 6b).
@@
We now demonstrate the emergence of anomalous nonlinear pumping in the continuous model in the presence of nonlinearity. The dynamics is governed by the following dimensionless Gross-Pitaevskii (GP) equation:
@@
where $\psi=\left(\psi_{1}(x),\psi_{2}(x)\right)^{T}$ and $\tilde{H}_{c}^{\textrm{lin}}$ is the dimensionless version of ${H}_{c}^{\textrm{lin}}$. The units of energy, time and length are $E_{R}$, $\hbar/E_{R}$ and $a$, respectively. We have taken $g_{11}=g_{22}=g$ for simplicity and have confirmed that the results remain valid when they are slightly different. We set $g=1$ and g=1g=-1 for repulsive and attractive interactions, respectively. The strength of nonlinearity is characterized by the norm $N=\sum_{\sigma}\int dx|\psi_{\sigma}(x)|^{2}$. For instance, consider a $^{7}\text{Li}$ BEC containing $1316$ atoms with the scattering length $a_{s}\approx-1.43$ nm under optical lattices along $x$ with $a=532$ nm and a transverse harmonic trap along $y$ and $z$ with the frequency $\omega_{\perp}=2\pi\times 710$ Hz [63, 64]. In this case, we have $N=0.2$.
By solving the nonlinear eigenstates of the instantaneous nonlinear Hamiltonian, we find that when $g>0$, the displacement of a soliton bifurcating from the first band over one period is $a$, consistent with linear Thouless pumping. However, the corresponding tight-binding model produces a nonlinear pumping displacement of $2a$, suggesting that the tight-binding model may oversimplify the nonlinear effects (see the discussion in Supplementary Information Section S-6). Remarkably, we find that when $g<0$, a soliton that bifurcates from the second band in the continuous model is pumped across 2a-2a over one cycle, as depicted in Fig. 6, which is twice the Chern number of the second band. We also find that for negative $g$, a soliton coming from the first band (whose Chern number is $1$) does not exhibit any pumping over one cycle (see Supplementary Information Section S-1), corresponding to the first anomalous case shown in Fig. 1a and Fig. 2b. Furthermore, we have numerically verified the stability of the instantaneous soliton solutions as $\theta$ runs from 0 to $2\pi$. In addition, the numerical time evolution of an initial soliton is also found to be in agreement with the instantaneous solution. These results suggest that anomalous nonlinear pumping can be experimentally observed in the continuous model.
Figure 6: Anomalous nonlinear pumping in the continuous model. a, One-cycle evolution of the density distribution $|\psi(x)|^{2}=\sum_{\sigma}|\psi_{\sigma}(x)|^{2}$ of the instantaneous solitons bifurcating from the second band when $g_{12}=0$. Note that the anomalous nonlinear pumping can still occur when $g_{12}\neq 0$. b, The evolution of center-of-mass positions $x_{c}$ of the instantaneous solitons (red line) and the corresponding Wannier functions (black line) with respect to $\theta$ over one period. Here, N=0.2{N}=0.2 and g=1g=-1.
@@
Discussion.—We have demonstrated the emergence of anomalous nonlinear Thouless pumping, wherein the correspondence between the displacement of a soliton and the Chern number of the underlying linear Bloch band breaks down. This anomalous behavior surpasses the understanding of nonlinear Thouless pumping based on the center-of-mass flow of the instantaneous Wannier functions. We show that the breakdown arises due to the transition of a soliton between different Wannier functions by passing through an intersite-soliton state. Furthermore, we illustrate that nonlinearity can induce the Thouless pumping of a soliton, even when the corresponding linear Bloch band is topologically trivial. We note that our results are completely different from the phenomenon in a Bose-Bose atomic mixture in Ref. [38] where a soliton of impurity atoms can undergo a nonzero quantized displacement during a pump cycle, as a result of the dragging from a soliton of majority atoms that bifurcates from a topologically nontrivial band. There, the pumping can also be understood as the center-of-mass flow of the Wannier functions of the corresponding Bloch band of the linear Hamiltonian felt by the majority atoms. In our case, we have only one set of band structures, and the discrepancy between the band topology and quantized motion emerges spontaneously, going beyond the description of the Wannier function of the original linear Hamiltonian. Given that anomalous nonlinear Thouless pumping can occur in both discrete and continuous models, it can be experimentally observed in either photonic systems or ultracold atomic gases. Additionally, our work inspires the exploration of other interesting topics, such as anomalous fractional Thouless pumping of solitons.